The following data were taken from the balance sheet of Nilo Company at the end of two recent fiscal years:

Current Year Previous Year
Current assets:
  Cash $396,700 $320,000
  Marketable securities 459,400 360,000
  Accounts and notes receivable (net) 187,900 120,000
  Inventories 382,800 244,000
  Prepaid expenses 197,200 156,000
  Total current assets $1,624,000 $1,200,000
Current liabilities:
  Accounts and notes payable
  (short-term) $336,400 $350,000
  Accrued liabilities 243,600 150,000
  Total current liabilities $580,000 $500,000

a. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round ratios to one decimal place.

Current Year Previous Year
1. Working capital $ $
2. Current ratio
3. Quick ratio

b. The liquidity of Nilo has   from the preceding year to the current year. The working capital, current ratio, and quick ratio have all  . Most of these changes are the result of an   in current assets relative to current liabilities.
Answer #1

a. Determine for each year (1) the working capital, (2) the current ratio, and (3) the quick ratio. Round ratios to one decimal place.

Current Year Previous Year
1. Working capital $1624000-580000 = 1044000 $1200000-500000 = 700000
2. Current ratio 1624000/580000 = 2.8 1200000/500000 = 2.4
3. Quick ratio 1044000/580000 = 1.8 800000/500000 = 1.6

b. The liquidity of Nilo has improved from the preceding year to the current year. The working capital, current ratio, and quick ratio have all liquidity ratio . Most of these changes are the result of an Increase in current assets relative to current liabilities.